Find a solution to dy/dx=xy+9x+4y+36




If necessary, use K to denote an arbitrary constant.


1
Expert's answer
2022-03-14T18:24:42-0400

The given equation can be written as, "\dfrac{dy}{dx} = (x+4)(y+9)"


Separating the variables, "\dfrac{dy}{y+9} = (x+4)dx"


Integrating we get,


"\begin{aligned}
\log (y+9) &= \dfrac{x^{2}}{2}+4x + c\\
y+9 &= e^{\frac{1}{2}(x^2+8x+2c)}\\
y&= Ke^{\frac{1}{2}(x^2+8x)} -9,~~~\text{where } K = e^{c}
\end{aligned}"